x - 3y + 6z = 21

3x + 2y - 5z = - 30

2x - 5y + 2z = -6

The sum of values of x

x, y

y, and z

z of the solution is


1
Expert's answer
2022-04-18T01:48:24-0400

Let's solve the system of linear equations using the Cramer's rule.

"\Delta=\begin{vmatrix}
 1& - 3&6\\
 3 & 2&-5\\
2&-5&2
\end{vmatrix}=\\
=1\cdot2\cdot2+(-3)\cdot(-5)\cdot2+6\cdot3\cdot(-5)-\\-6\cdot2\cdot2-1\cdot(-5)\cdot(-5)-(-3)\cdot3\cdot2=\\
=4+30-90-24-25+18=-87;"


"\Delta_1=\begin{vmatrix}
 21& - 3&6\\
 -30 & 2&-5\\
-6&-5&2
\end{vmatrix}=\\
=21\cdot2\cdot2+(-3)\cdot(-5)\cdot(-6)+6\cdot(-30)\cdot(-5)-\\-6\cdot2\cdot(-6)-21\cdot(-5)\cdot(-5)-(-3)\cdot(-30)\cdot2=\\
=84-90+900+72-525-180=261;"


"\Delta_2=\begin{vmatrix}
 1& 21&6\\
 3 & - 30&-5\\
2&-6&2
\end{vmatrix}=\\
=1\cdot(-30)\cdot2+21\cdot(-5)\cdot2+6\cdot3\cdot(-6)-\\-6\cdot(-30)\cdot2-1\cdot(-5)\cdot(-6)-21\cdot3\cdot2=\\
=-60-210-108+360-30-126=-174;"


"\Delta_3=\begin{vmatrix}
 1& - 3&21\\
 3 & 2&-30\\
2&-5&-6
\end{vmatrix}=\\
=1\cdot2\cdot(-6)+(-3)\cdot(-30)\cdot2+21\cdot3\cdot(-5)-\\-21\cdot2\cdot2-1\cdot(-30)\cdot(-5)-(-3)\cdot3\cdot(-6)=\\
=-12+180-315-84-150-54=-435;"


"x=\cfrac{\Delta_1} {\Delta} =\cfrac{261}{-87}-3;\\
y=\cfrac{\Delta_2} {\Delta} =\cfrac{-174}{-87}=2;\\
z=\cfrac{\Delta_3} {\Delta} =\cfrac{-435}{-87}=5;\\
x+y+z=-3+2+5=4."